FOURIER TRANSFORM MICROWAVE SPECTRUM OF THE FeCN RADICAL (X^{4}Δ_{i}) AND CONFIRMA-TION OF THE GROUND ELECTRONIC STATE

\underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721; M. A. FLORY, CNA, Frankfort, KY; B. J. HARRIS, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the FeCN radical (X^{4}Δ_{i}) has been measured using Fourier transform microwave (FTMW) techniques in the 4 - 40 GHz frequency range and a laser ablation source. The species was produced using Discharge Assisted Laser Ablation Spectroscopy (DALAS) in a supersonic jet expansion of iron vapor and (CN)_{2}, diluted in argon carrier gas. The fundamental rotational transition, J = 9/2 → 7/2 (Ω = 7/2), was recorded near 36 GHz. Three hyperfine transitions, due to the nitrogen nuclear spin of I = 1, were observed in the spectrum. These data were combined with the previous millimeter/submillimeter measurements of FeCN in a global fit and nitrogen hyperfine constants were determined. These measurements confirm that the ground state of FeCN is ^{4}Δ_{i}, as suggested by previous millimeter/submillimeter measurements of Flory & Ziurys. Theoretical calculations have predicted that the ground state of this radical is ^{6}Δ_{i}. In a ^{6}Δ_{i} state, the J = 9/2 → 7/2 transition does not exist in the lowest energy ladder (Ω = 9/2).